OpenStudy (anonymous):

write an equation in piont-slope form if the line that passes through the given piont and with the given slope m. (3 -4);m=6 im confuzzed what do i do ?

OpenStudy (turingtest):

point-slope form is\[y-y_1=m(x-x_1)\]you know what m is, and \[(x_1,y_1)\]is the point given in the problem, so just plug in the numbers.

OpenStudy (anonymous):

My answer: y + 4 = 6(x - 3) ------ The equation.

OpenStudy (turingtest):

and she's right

OpenStudy (anonymous):

i still do not get how to do it

OpenStudy (turingtest):

your point is\[(x_1,y_1)=(3,-4)\]and you have \[m=6\]plug those numbers into\[y-y_1=m(x-x_1)\]and you're done